OpenStudy (anonymous):

What is the equation of the line that passes through (1, 3) and (-2, -3)?

OpenStudy (anonymous):

First, let's find the slope. Slope = rise / run (or I always remember it that way), so the difference in Y over the difference in X. (3-(-3)) / (1-(-2)) = 6/3 = 2. Now, we know y = 2x + b but need to find b. Thankfully, we have a point. 3 = 2(1) + b b = 1 y = 2x + 1

OpenStudy (anonymous):

thanks !